OpenStudy (anonymous):

Rewrite the expressions without using absolute value notation |x+1| +4 |x+3| given that -(5/2)

OpenStudy (anonymous):

\[\frac{-5}{2}<x<\frac{-3}{2}\] \[\frac{-5}{2}+1<x+1<\frac{-3}{2}+1\] \[\frac{-3}{2}<x+1<\frac{-1}{2}\] x+1 is <0 in the interval therefore |x+1|=-(x+1) x+3 is >0 in the interval |x+3|=x+3